# Who Are We Missing?: A Principled Approach to Characterizing the Underrepresented Population

### v2

@@ 1 Introduction
Randomized controlled trials (RCTs) are the gold standard for estimating the causal effects of interventions and informing decision-making (Ding,, 2023). However, a challenge can lie in extending the inferences from trials to target populations, particularly in the face of effect heterogeneity (Degtiar and Rose,, 2023; Rudolph et al.,, 2023). Prior research underscores that the generalizability of RCT evidence to the target population depends on the alignment of distributions of individuals’ characteristics - in particular, effect moderators - between the trial cohort and the target population relevant for decision-making. Differences in distributions can compromise treatment effect extrapolations, especially for groups underrepresented in trials, potentially leading to inaccuracies and incorrect decision-making (Bell et al.,, 2016; Olsen et al.,, 2013; Imai et al.,, 2008; Pearl and Bareinboim,, 2011, 2014; Bareinboim et al.,, 2014). It can be particularly hard to identify when such misalignment exists in high-dimensional settings. Crucially, comparability primarily matters for moderators (also known as causal effect modifiers), but identifying which characteristics are moderators can be challenging.
Concerns about the potential lack of representativeness of trial participants have been raised in studies of emergency contraception (Glasier et al.,, 2011), education (Tipton and Olsen,, 2018), substance use (Rudolph et al.,, 2022), etc. For example, Glasier et al., (2011) demonstrated that the risk of becoming pregnant for obese women was over threefold higher compared to women with a normal body mass index (BMI). Yet, individuals with a high BMI were frequently underrepresented in emergency contraception clinical trials. Consequently, decisions informed by these trials placed them at an unwarranted risk of pregnancy, highlighting the real-world implications of underrepresentation in clinical trials. Similarly, recent literature underscores the challenges posed by underrepresentation and calls for trials that more accurately represent diverse populations (Boden-Albala,, 2022; Bierer et al.,, 2022; Rudolph et al.,, 2022). However, as discussed earlier, in practical terms, determining which features to prioritize and identifying underrepresented subgroups remains a complex challenge.
Contribution.
Our work proposes a principled approach to identify underrepresented groups for whom a trial yields insufficient information regarding treatment effect estimates. Once identified, the target population can be refined by removing underrepresented groups, ensuring a more precise estimate of treatment effects and clarity regarding the groups to whom generalization is – or isn’t – warranted. Our framework translates the objective to an optimization problem that learns the target subpopulation distribution that minimizes the variance of the target average treatment effect (TATE) estimate. We introduce a nonparametric functional optimization approach, Rashomon Set of Optimal Trees (ROOT), to solve the optimization problem. We ensure the output of ROOT is interpretable, enabling researchers to effectively communicate the characteristics of the study population as well as the underrepresented groups. This can inform where the trial inferences can be most precisely extended as well as alert those who may use the trial study results to particular subpopulations underrepresented in the current study. We rigorously evaluated our approach through synthetic data experiments involving complex data-generative procedures with non-linear relationships. We compared the performance of our optimization approach against simpler alternatives. Our findings demonstrate that our approach is efficient in improving precision and yields interpretable results.
@@
We apply our approach to 1) identify underrepresented groups in an RCT of medication for opioid use disorder (MOUD), and 2) extend inferences from that trial to a refined target population that removes these underrepresented groups. This motivating example is of interest, because, although MOUD (methadone, buprenorphine, and extended-release naltrexone) are effective in reducing the risk of opioid overdose, their effect sizes estimated in trials are generally appreciably larger in trials than observed in real-world clinical scenarios (Rudolph et al.,, 2022; Susukida et al.,, 2016; Blanco et al.,, 2017), and it is important to understand why. We hypothesize that a portion of this incongruity stems from differences between the individuals in the trials and the broader population of patients treated for opioid use disorder. Specifically, we consider the Starting Treatment with Agonist Replacement Therapies (START) MOUD trial (Saxon et al.,, 2013), which compares methadone with buprenorphine; and the real-world population of people for whom MOUD is part of their treatment plan, using the Treatment Episode Dataset: Admissions (TEDS-A) (Abuse et al.,, 2020), which includes all people treated for substance use disorders at programs across the United States that receive public funding.
@@ 3 Preliminaries
In this study, we are interested in generalizing the causal effect of treatment ($T$) on outcome ($Y$) from an RCT to a target population. Our dataset has $n$ individuals $\mathcal{D}_{n}={1,\ldots,n}$. Corresponding to each unit $i\in\mathcal{D}_{n}$ and binary treatment {1,0}, there are two potential outcomes, represented as $Y_{i}(1)$ and $Y_{i}(0)$, respectively. We define the individual treatment effect, $\tau_{i}$, as the difference between these potential outcomes for person $i$, $\tau_{i}:=Y_{i}(1)-Y_{i}(0)$. Each unit $i\in\mathcal{D}_{n}$ is categorized as either belonging to the RCT ($S_{i}=1$) or the target population ($S_{i}=0$). In the RCT, we observe a $p$-dimensional vector of pretreatment covariates ($\mathbf{X}_{i}=(X_{1,i},\ldots,X_{p,i}$)), a treatment indicator ($T_{i}$), and one of the two potential outcomes for each person, depending on their treatment assignment: $Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(0)$. However, for units in the target population ($S_{i}=0$), we observe only their covariates $\mathbf{X}_{i}$. To simplify our discussion, we assume Bernoulli randomization in the RCT. We denote $n_{1}$ as the number of units in the RCT ($S=1$) and $n_{0}$ as the number of units in the target sample ($S=0$), such that $n=n_{1}+n_{0}$. Our primary estimands of interest are the target average treatment effect (TATE): ${\tau}_{0}:=\mathbb{E}[\tau_{i}|S_{i}=0]$, and the sample target average treatment effect (STATE): ${\tau}_{0}^{\mathcal{D}_{n}}:=\frac{1}{n_{0}}\sum_{i}(1-S_{i})\tau_{i}$. Here, STATE is a finite sample equivalent version of the TATE.
@@
    (Internal Validity) In the RCT, treatment assignment is independent of the potential outcomes conditional on pre-treatment covariates X: $(Y(1),Y(0))\perp T\mid\mathbf{X},S=1$.
@@
Assumption A.1. establishes the support of the population distribution, ensuring that the probability of observing any point within the support is non-zero. This standard assumption is fundamental in causal inference, as it guarantees the existence of a comparable unit to estimate the missing counterfactual. The second part of the assumption, $10<P(T=1\mid\mathbf{X},S=1)<$ 1, is guaranteed due to the design of the experiment; people enrolled in a trial must be able to be randomized to either treatment. However, the first part, $10<P(S=1\mid\mathbf{X})<$ 1, may not hold, e.g., when units with specific covariate values are systematically excluded from the RCT or never receive treatment. Assumption A.2. is generally upheld in well-implemented trials. Assumption A.3. implies that there are no unobserved factors differentiating units in the trial from the population. Under this assumption, $P(Y(t)|\mathbf{X}=\mathbf{x},S=1)=P(Y(t)|\mathbf{X}=\mathbf{x},S=0)$.
@@
We define a few more quantities for notational convenience and parsimony: (i) expected outcome, $\mu(\mathbf{X},t):=\mathbb{E}(Y\mid\mathbf{X},T=t,S=1)$, (ii) selection score, $\pi(\mathbf{X}):=P(S=1\mid\mathbf{X})$, (iii) selection odds, $\ell(\mathbf{X}):=\frac{\pi(\mathbf{X})}{1-\pi(\mathbf{X})}$, (iv)$\pi:=P(S=1)$, and (v) trial propensity score, $e(\mathbf{X}):=P(T=1\mid\mathbf{X},S=1)$. Then, given assumptions A.1.A.3., TATE ${\tau}_{0}$ is identified by $\theta_{0}:=\mathbb{E}_{\mathbf{X}\mid S=0}(\mu(\mathbf{X},{1})-\mu(\mathbf{X},{0}))\mathrm{,and}$ $\theta_{1}:=\frac{\pi}{1-\pi}\mathbb{E}_{\mathbf{X}\mid S=1}\left(\frac{\mu(\mathbf{X},{1})-\mu(\mathbf{X},{0})}{\ell(\mathbf{X})}\right).$ These results are well-known in the literature and shown in Rudolph and Laan, (2017) and Rudolph et al., (2023). Proof of this result as well as all the subsequent results are in Appendix A. Next, we focus on the estimation of TATE.
@@
One of the estimators for TATE that is inspired by $\theta_{1}$ is the inverse probability-weighted (IPW) estimator. We focus on this estimator for simplicity but note that other estimators exist, including an outcome model-based approach and a doubly robust approach combining outcome model and IPW approaches. Let $\widehat{Y}^{ipw}_{i}(t)=\frac{\widehat{\pi}}{1-\widehat{\pi}}\frac{S_{i}\mathbbm{1}(T_{i}=t)Y_{i}}{\widehat{\ell}(\mathbf{X}_{i})\widehat{e}(\mathbf{X}_{i})}$. Then
@@
$\mathbb{V}^{ipw}:=\left(\frac{\pi}{1-\pi}\right)^{2}\mathbb{E}_{\mathbf{X}\mid S=1}\left[\frac{\begin{pmatrix}\sigma^{2}(\mathbf{X},1)\frac{1-e(\mathbf{X})}{e(\mathbf{X})}+\sigma^{2}(\mathbf{X},0)\frac{e(\mathbf{X})}{1-e(\mathbf{X})}\\ +\left(\mu(\mathbf{X},1)\left(\frac{1-e(\mathbf{X})}{e(\mathbf{X})}\right)^{1/2}+\mu(\mathbf{X},0)\left(\frac{e(\mathbf{X})}{1-e(\mathbf{X})}\right)^{1/2}\right)^{2}\end{pmatrix}}{(\ell(\mathbf{X}))^{2}}\right].$
@@
An essential consideration in generalizing trial estimates lies in the precision of the effect estimates. As seen in Equation 1, the precision of estimates diminishes for the subspace with smaller $\ell(\mathbf{X})$. A potential strategy for ensuring a precise estimate thus involves constraining the analysis to a subspace with larger $\ell(\mathbf{X})$. While this may enhance the precision of estimates, the drawback is the reduction in generalizability as well as the sample size $n_{1}$, as units with low $\ell(\mathbf{X})$ are pruned. This reduction in sample size can decrease the precision of the TATE estimate.
@@
We now consider the numerator inside the expectation:
$\sigma^{2}(\mathbf{X},1)\frac{1-e(\mathbf{X})}{e(\mathbf{X})}+\sigma^{2}(\mathbf{X},0)\frac{e(\mathbf{X})}{1-e(\mathbf{X})}+\left(\mu(\mathbf{X},1)\left(\frac{1-e(\mathbf{X})}{e(\mathbf{X})}\right)^{1/2}+\mu(\mathbf{X},0)\left(\frac{e(\mathbf{X})}{1-e(\mathbf{X})}\right)^{1/2}\right)^{2}.$
The first two terms quantify the stochasticity or noise in the outcome from the trial, while the third term gauges the scale of the conditional outcomes. It is important to note that while none of these terms explicitly address precision loss due to underrepresentation, they indicate that precise estimation of the TATE requires distributional alignment on covariates that are effect modifiers.
@@ 4 Shifting the Goalpost
${\tau}_{0}^{w}=\frac{\mathbb{E}[w(\mathbf{X})\tau(\mathbf{X})\mid S=0]}{\mathbb{E}[w(\mathbf{X})\mid S=0]}=\frac{\mathbb{E}[w(\mathbf{X})\tau(\mathbf{X})/\ell(\mathbf{X})\mid S=1]}{\mathbb{E}[w(\mathbf{X})/\ell(\mathbf{X})\mid S=1]}.$ (2)
Here, a choice of binary $w(\mathbf{x})$ corresponds to including or excluding the individuals with $\mathbf{X}$ equal to $\mathbf{x}$ in the study population. Analogously, it also corresponds to whether the individuals with covariate values $\mathbf{x}$ are represented or underrepresented in the study population.
@@
$\widehat{Y}^{ipw}_{i}(t)$ $=\frac{\widehat{\pi}}{1-\widehat{\pi}}\frac{S_{i}\mathbbm{1}(T_{i}=t)Y_{i}}{\widehat{\ell}(\mathbf{X}_{i})\widehat{e}(\mathbf{X}_{i},t)}$
@@
where
${\small\mathbb{V}^{ipw,w}:=\left(\frac{\pi/(1-\pi)}{\mathbb{E}_{\mathbf{X}|S=1}\left[w(\mathbf{X})\right]}\right)^{2}\mathbb{E}_{\mathbf{X}\mid S=1}\left[\left(\frac{w(\mathbf{X})}{\ell(\mathbf{X})}\right)^{2}\alpha^{2}(\mathbf{X})\right]},$
and $\alpha^{2}(\mathbf{X})=\sigma^{2}(\mathbf{X},1)\frac{1-e(\mathbf{X})}{e(\mathbf{X})}+\sigma^{2}(\mathbf{X},0)\frac{e(\mathbf{X})}{1-e(\mathbf{X})}+\left(\mu(\mathbf{X},1)\left(\frac{1-e(\mathbf{X})}{e(\mathbf{X})}\right)^{1/2}+\mu(\mathbf{X},0)\left(\frac{e(\mathbf{X})}{1-e(\mathbf{X})}\right)^{1/2}\right)^{2}.$ Here, the problem of characterizing the target population where we can precisely estimate the TATE reduces to finding optimal $w$ such that
@@
$\widehat{\mathbb{V}}^{ipw,w}_{\mathcal{D}_{n}}=\frac{n_{1}}{(\sum_{i}S_{i}w(\mathbf{X}_{i}))^{2}}\sum_{i}w(\mathbf{X}_{i})S_{i}\left(\widehat{Y}^{ipw}_{i}(1)-\widehat{Y}^{ipw}_{i}(0)-\widehat{\tau}_{0}^{ipw,w}\right)^{2}.$
@@
Not only do we want to learn an optimal function $w$ using the data $\mathcal{D}_{n}$ but also we want $w$ to be “interpretable” to ensure that we can decribe who is and is not included in the study population. Modeling $w$ as a function of summaries produced by $\ell$ and $\mu$ would not yield the interpretability needed. Instead, we model $w$ directly as a function of the covariates, $\mathbf{X}$.

### v4

@@ Abstract
Keywords: Causal Inference, Generalizability, Data Fusion, Machine Learning, Interpretability
@@ 1 Introduction
Randomized controlled trials (RCTs) are the gold standard for estimating treatment effects due to the internal validity they achieve by randomly assigning participants to treatment arms (Ding,, 2023). However, the characteristics of the trial cohort often differ from the target population about whom one would like to draw inferences (Degtiar and Rose,, 2023). Further, if the misaligned population characteristics between the trial and the target population are also effect modifiers, this can challenge the generalizability of trial findings (Degtiar and Rose,, 2023; Rudolph et al., 2023b, ).
Existing approaches generalizing trial inferences adjust for differences in the distribution of population characteristics between the trial cohort and the target population (Degtiar and Rose,, 2023). However, many of these estimators lack precision while estimating effects over subgroups insufficiently represented in the trial cohort. This can lead to treatment decisions based on limited data and strong assumptions. Despite this concern, there is limited focus on identifying regions of the target population where inferences are most reliable and, crucially, areas where the trial provides little reliable information. Existing methods that focus on characterizing underrepresented subgroups do not account for outcome variability and only account for sample selection probability (Stuart et al.,, 2011; Tipton,, 2014; Ding et al.,, 2016; Tipton and Peck,, 2017; Cahan et al.,, 2017; Greenhouse et al., 2008a, ). Further, they rely on summary measures such as selection scores that are inherently hard to interpret.
Contributions.
In this paper, we aim to interpretably characterize subgroups for which precise estimation of treatment effects is particularly challenging. These subgroups often occupy regions of the covariate space with heterogeneous treatment effects and insufficient representation in the trial data. We refer to these subgroups as “insufficiently represented” or “underrepresented” populations, due to their lack of data support in the trial, which hampers precise treatment effect estimation. Our approach involves identifying trial units that significantly contribute to the variance in estimating the target average treatment effect (TATE). We also redefine the estimand to focus on inferences within sufficiently represented subpopulations. This introduces two tensions: (i) a conceptual tension between confidence in the inferences versus the extent of their generalizability, and (ii) a computational tension between dropping units to minimize variance versus maximizing sample size for statistical power.
We formalize this computational tension as a functional optimization problem and introduce an interpretable, non-parametric approach for functional optimization called the Rashomon Set of Optimal Trees (ROOT) [^1] Rashomon is a classic Japanese movie in which an event is given more than one and often a competing interpretation by the individuals involved. It provides different and sometimes contradictory points of view on the same event. Here, each of the trees in the set provides almost equally good but the different characterization of underrepresented subpopulations – thus we refer to this set as a Rashomon Set. Analogously, there is also literature in Interpretable Machine Learning that focuses on understanding the properties and behavior of almost equally good but different models (Breiman,, 2001; Semenova et al.,, 2022).. By interpretable, we mean that researchers can effectively communicate the characteristics of insufficiently represented groups – this helps address the conceptual tension by characterizing the population for whom the inferences can be made with high confidence. Given the trial and target data, ROOT identifies trial subpopulations that contribute significantly to the variance of the target treatment effect estimate. This high variance arises from a lack of balance in potential effect modifiers between the trial and target samples. In other words, ROOT identifies subpopulations with a small estimated probability of being in the trial population conditional on the effect modifiers—i.e., those units that may exhibit practical violations of positivity (discussed technically in Section 3). ROOT uses a tree-sampling-based algorithm with an explore-exploit search strategy to find the optimal solution to the functional optimization problem, accounting for potential effect heterogeneity and differences in characteristics between trial and target populations. We construct a set of trees with near-optimal objective values, recognizing the existence of multiple equally good models with different representations and interpretations, called a Rashomon set (Xin et al.,, 2022).
@@
We apply our approach to investigate the research-to-practice gap in studies of medication for opioid use disorder (MOUD), where discrepancies between treatment effectiveness measured in trials and real-world care settings have been noted. Recent work has highlighted that this discrepancy can be due to underrepresentation in trial cohorts (Rudolph et al.,, 2022). We address this by identifying underrepresented groups in an RCT of MOUD and extending inferences from that trial to a refined target population that excludes these underrepresented groups. Specifically, we consider the Starting Treatment with Agonist Replacement Therapies (START) MOUD trial (Saxon et al.,, 2013; Hser et al.,, 2014), which compares methadone with buprenorphine, and the real-world population of individuals for whom MOUD is part of their treatment plan, using the Treatment Episode Dataset: Admissions (TEDSA) (Abuse et al.,, 2020). This dataset includes individuals treated for substance use disorders at programs across the United States that receive public funding.
Paper Organization.
We first summarize the relevant literature in Section 2. In Section 3, we introduce our setup and assumptions. We propose our framework and the ROOT approach in Sections 4 and 5. We then proceed to demonstrate the effectiveness of our approach through synthetic data experiments in Section D. Lastly, in Section 6.2, we discuss the application of our methodology to the MOUD case study. We end with a discussion of the limitations of our approach, posing open questions and future directions (Section 7).
@@ 3 Preliminaries
In this study, we are interested in generalizing the causal effect of treatment ($T$) on outcome ($Y$) from an RCT to a target population. Our dataset comprises of $n$ individuals $\mathcal{D}_{n}={1,\ldots,n}$. Corresponding to each unit $i\in\mathcal{D}_{n}$ and binary treatment choice, $1$ and 00, there are two potential outcomes, represented as $Y_{i}(1)$ and $Y_{i}(0)$, respectively. We define the individual treatment effect, $\tau_{i}$, as the difference between these potential outcomes for person $i$, $\tau_{i}:=Y_{i}(1)-Y_{i}(0)$. Each unit $i\in\mathcal{D}_{n}$ is categorized as either belonging to the RCT ($S_{i}=1$) or the target population data ($S_{i}=0$). In the RCT, we observe a $p$-dimensional vector of pretreatment covariates ($\mathbf{X}_{i}=(X_{1,i},\ldots,X_{p,i}$)), a treatment indicator ($T_{i}$), and one of the two potential outcomes for each person, depending on their treatment assignment: $Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(0)$. However, for units in the target population ($S_{i}=0$), we observe only their covariates $\mathbf{X}_{i}$. To simplify our discussion, we assume Bernoulli randomization in the RCT. For clarity, we denote $n_{1}$ as the number of units in the RCT ($S=1$) and $n_{0}$ as the number of units in the target sample ($S=0$), such that $n=n_{1}+n_{0}$. Our primary estimands of interest are the target average treatment effect (TATE): ${\tau}_{0}:=\mathbb{E}[\tau_{i}|S_{i}=0]$, and the sample target average treatment effect (STATE): ${\tau}_{0}^{\mathcal{D}_{n}}:=\frac{1}{n_{0}}\sum_{i}(1-S_{i})\tau_{i}$. Here, STATE is a finite sample equivalent version of the TATE.
@@
    (Internal Validity) In the RCT, treatment assignment is independent of the potential outcomes: $(Y(1),Y(0))\perp T\mid\mathbf{X},S=1$.
@@
  4. A.4.
    (Stable Unit Treatment Value) For unit $i$ in the RCT, i.e. $S_{i}=1$ and any pair of treatment assignments $\{T_{1},\dots,T_{n_{1}}\}$ and $\{T^{\prime}_{1},\dots,T^{\prime}_{n_{1}}\}$, if $T_{i}=T^{\prime}_{i}$ then $Y_{i}(T_{i})=Y_{i}(T^{\prime}_{i})$. Further, for any pair of sample assignments $\{S_{1},\dots,S_{n}\}$ and $\{S^{\prime}_{1},\dots,S^{\prime}_{n}\}$, if $S_{i}=S^{\prime}_{i}$ then $Y_{i}(t,S_{i})=Y_{i}(t,S^{\prime}_{i})$ for all $t\in\{0,1\}$.
@@
These assumptions are also discussed in Stuart et al., (2011) and Tipton, (2013). Assumption A.1. establishes the support of the population distribution, ensuring that the probability of observing any point within the support is non-zero. This standard assumption is fundamental in causal inference, as it guarantees the existence of a comparable unit to estimate the missing counterfactual. While the second part of the assumption – $10<P(T=1\mid\mathbf{X},S=1)<$ 1 – is guaranteed due to the design of the experiment; people are not enrolled in a trial if they can’t have either treatment. However, the first part – $10<P(S=1\mid\mathbf{X})<$ 1 – can be challenged when units with specific covariate values are systematically excluded from the RCT or never receive treatment. In our work, we assume that structural positivity holds i.e., $P(S=1\mid\mathbf{X}=\mathbf{x})\neq 0$, however, we focus on scenarios with practical violations of positivity (as defined in Petersen et al., (2012)) where certain subpopulations are insufficiently represented in the trial cohort. Assumption A.2. is generally upheld in well-implemented trials. Assumption A.3. implies that there are no unobserved factors differentiating units in the trial from the population. Under this assumption, $P(Y(t)|\mathbf{X}=\mathbf{x},S=1)=P(Y(t)|\mathbf{X}=\mathbf{x},S=0)$. This assumption is also referred to as “S-admissibility” (Egami and Hartman,, 2023; Pearl and Bareinboim,, 2011).
@@
We define a few more quantities for notational convenience and parsimony: (i) expected outcome, $\mu(\mathbf{X},t):=\mathbb{E}(Y\mid\mathbf{X},T=t,S=1)$, (ii) selection score, $\pi(\mathbf{X}):=P(S=1\mid\mathbf{X})$, (iii) selection ratio, $\ell(\mathbf{X}):=\frac{\pi(\mathbf{X})}{1-\pi(\mathbf{X})}$, (iv)$\pi:=P(S=1)$, and (v) trial propensity score, $e(\mathbf{X}):=P(T=1\mid\mathbf{X},S=1)$. Then, given assumptions A.1.A.3., TATE ${\tau}_{0}$ is identified by $\theta_{0}:=\mathbb{E}_{\mathbf{X}\mid S=0}(\mu(\mathbf{X},{1})-\mu(\mathbf{X},{0}))\mathrm{,and}$ $\theta_{1}:=\frac{\pi}{1-\pi}\mathbb{E}_{\mathbf{X}\mid S=1}\left(\frac{\mu(\mathbf{X},{1})-\mu(\mathbf{X},{0})}{\ell(\mathbf{X})}\right).$ These results are well-known in the literature and shown in Rudolph and Laan, (2017) and Rudolph et al., 2023b . Proof of this result as well as all the subsequent results are in Appendix A. Next, we focus on the estimation of TATE.
@@
For ease of illustration, we introduce our approach for a simple inverse probability-weighted (IPW) estimator for the TATE proposed in Stuart et al., (2011). A similar and analogous analysis can be carried out using any consistent estimator of TATE including an outcome model-based approach and a doubly robust approach combining outcome model and IPW approaches. Let $\widehat{Y}^{ipw}_{i}(t)=\frac{\widehat{\pi}}{1-\widehat{\pi}}\frac{S_{i}\mathbbm{1}(T_{i}=t)Y_{i}}{\widehat{\ell}(\mathbf{X}_{i})\widehat{e}(\mathbf{X}_{i})}$. Then
@@
$\mathbb{V}^{ipw}=\frac{\pi}{(1-\pi)^{2}}\mathbb{E}_{\mathbf{X}|S=1}\left[\left(\frac{\sigma^{2}(\mathbf{X},1)}{\ell^{2}(\mathbf{X})e(\mathbf{X})}+\frac{\sigma^{2}(\mathbf{X},0)}{\ell^{2}(\mathbf{X})(1-e(\mathbf{X}))}\right)+\frac{(1-\pi)^{2}}{\pi(\mathbf{X})}\left({\tau}_{0}(\mathbf{X})-{\tau}_{0}\right)^{2}\right].$ (2)
@@
An essential consideration in generalizing trial estimates lies in the precision of the effect estimates. As seen in Equation 2, the contribution to the variance is larger for smaller $\pi(\mathbf{x})$ i.e., the precision of estimates diminishes for the subspace exhibiting “practical violations of the positivity” (Petersen et al.,, 2012). A potential strategy for ensuring a precise estimate thus involves constraining the analysis to a subspace with larger $\pi(\mathbf{x})$. While this may enhance the precision of estimates, the drawback is the reduction in generalizability as well as the sample size $n_{1}$, as units with low $\pi(\mathbf{x})$ are pruned. This reduction in sample size can decrease the precision of the TATE estimate.
@@
Now, let’s shift our focus to the second term: $(1-\pi)^{2}\frac{\left({\tau}_{0}(\mathbf{x})-{\tau}_{0}\right)^{2}}{\pi(\mathbf{x})}$. The numerator measures level of heterogeneity of the treatment effect i.e. how different is ${\tau}_{0}(\mathbf{x})$ from ${\tau}_{0}$, and the denominator $\pi(\mathbf{x})$, as before, measures how the representativeness of units with covariate $\mathbf{X}=\mathbf{x}$. If $\pi(\mathbf{x})$ is small, then it implies that there is limited overlap between the target and trial data in the region of covariate space with $\mathbf{X}=\mathbf{x}$. However, if ${\tau}_{0}(\mathbf{x})$ is not very different from ${\tau}_{0}$ then this limited overlap does not contribute high amount of variance or uncertainty in the estimation. This indicates that precise estimation of the TATE requires distributional alignment on covariates that are effect modifiers. This result is well known in the literature.(e.g., Rudolph et al., 2023b, ; Zeng et al.,, 2023)
@@ 4 Shifting the Goalpost
${\tau}_{0}^{w}=\frac{\pi_{w}}{1-\pi_{w}}\frac{\mathbb{E}[w(\mathbf{X}){\tau}_{0}(\mathbf{X})/\ell(\mathbf{X})\mid S=1]}{\mathbb{E}[w(\mathbf{X})\mid S=1]},$ (3)
where, $\pi_{w}=\mathbb{E}[w(\mathbf{X})S]$ and ${\tau}_{0}(\mathbf{X})=\mathbb{E}[Y(1)-Y(0)\mid\mathbf{X},S=0]$. Here, a choice of binary $w(\mathbf{x})$ corresponds to including or excluding the individuals with $\mathbf{X}$ equal to $\mathbf{x}$ in the study population. Analogously, it also corresponds to whether the individuals with covariate values $\mathbf{x}$ are represented or underrepresented in the study population.
@@
$\widehat{Y}^{ipw,w}_{i}(t)$ $=\frac{\widehat{\pi}_{w}}{1-\widehat{\pi}_{w}}\frac{S_{i}\mathbbm{1}(T_{i}=t)Y_{i}}{\widehat{\ell}(\mathbf{X}_{i})\widehat{e}(\mathbf{X}_{i},t)}$
@@
and $\alpha^{2}(\mathbf{X})=\left(\frac{\sigma^{2}(\mathbf{X},1)}{\ell^{2}(\mathbf{X})e(\mathbf{X})}+\frac{\sigma^{2}(\mathbf{X},0)}{\ell^{2}(\mathbf{X})(1-e(\mathbf{X}))}\right)+\frac{(1-\pi)^{2}}{\pi(\mathbf{X})}\left({\tau}_{0}(\mathbf{X})-{\tau}_{0}^{w}\right)^{2}.$ Here, the problem of characterizing the target population where we can precisely estimate the TATE reduces to finding optimal $w$ such that
@@
$\widehat{\mathbb{V}}^{ipw,w}_{\mathcal{D}_{n}}=\frac{n_{1}}{(\sum_{i}S_{i}w(\mathbf{X}_{i}))^{2}}\sum_{i}w^{2}(\mathbf{X}_{i})S_{i}\left(\widehat{Y}^{ipw,w}_{i}(1)-\widehat{Y}^{ipw,w}_{i}(0)-\widehat{\tau}_{0}^{ipw,w}\right)^{2}.$
@@
Not only do we want to learn an optimal function $w$ using the data $\mathcal{D}_{n}$ but also we want $w$ to be “interpretable” to ensure that we can decribe who is and is not included in the study population. Modeling $w$ as a function of summaries produced by $\ell$ and $\mu$ would not yield the interpretability needed. We model $w$ directly as a function of the covariates, $\mathbf{X}$.